OpenStudy (stuck-help):

METAL AND FAN ( for the right answer)Suppose you divide a polynomial by a binomial. How do you know if the binomial is a factor of the polynomial? Create a sample problem that has a binomial which IS a factor of the polynomial being divided, and another problem that has a binomial which is NOT a factor of the polynomial being divided.

OpenStudy (drbcaptin):

http://www.jiskha.com/search/index.cgi?query=Binomial+Theoram See if this helps you out.

OpenStudy (anonymous):

After you've done the division, you know the binomial is a factor if the remainder is zero. (3x^3 - 2x^2 - 2x - 5) / (3x - 5) = x^2 + x + 1 Since there is no remainder, 3x - 5 is a factor of 3x^3 - 2x^2 - 2x - 5. (3x^3 - 2x^2 - 2x - 7) / (3x - 5) = x^2 + x + 1 with remainder -2. Since the remainder is not zero, 3x - 5 is not a factor of 3x^3 - 2x^2 - 2x - 7.

OpenStudy (anonymous):

hey jprose orDRBCaptain can one of you guys help i posted a question

OpenStudy (anonymous):

sure